OpenStudy (anonymous):

factor 24a^3b^4-6a^5b^7-21ab^3 I really need help with factoring

OpenStudy (anonymous):

\[24a^3b^4-6a^5b^7-21ab^3\]

OpenStudy (anonymous):

yesss thats the equation

OpenStudy (anonymous):

alright then look at the terms and can you see anything that they all share?

OpenStudy (anonymous):

ab??

OpenStudy (anonymous):

ok so we pull ab out of the entire expression \[(ab)24a^2b^3-6a^4b^6-21b^2\] can you see anything else that they all share?

OpenStudy (anonymous):

oops i typed that wrong it should be

OpenStudy (anonymous):

\[(ab)(24a^2b^3-6a^4b^6-21b^2)\]

OpenStudy (anonymous):

they all share.....?? exponents???/

OpenStudy (anonymous):

ok just look at this expression \[24a^2b^3-6a^4b^6-21b^2\] ignore the ab for now and remember your times table for 3

OpenStudy (anonymous):

what else does each value in that expression share?

OpenStudy (anonymous):

they share multiples of 3

OpenStudy (anonymous):

ok what else?

OpenStudy (anonymous):

ummmm...? thats all i can think of...

OpenStudy (anonymous):

let me remind you that b^5 = b*b*b*b*b b^2= b*b

OpenStudy (anonymous):

ok i got that

OpenStudy (anonymous):

and b^5 = b^2*b^3

OpenStudy (anonymous):

ok then ill make it more obvious \[24a^2b^2b-6a^4b^2b^4-21b^2\]

OpenStudy (anonymous):

using the distribution property of multiplication, what else can we pull out of this expression?

OpenStudy (anonymous):

i don't get why you just did that though...

OpenStudy (anonymous):

they all have b^2 in common...

OpenStudy (anonymous):

they all have a b

OpenStudy (anonymous):

so we pulled out (ab)(3)(b^2) \[24a^3b^4-6a^5b^7-21ab^3=(3)(ab)(b^2)(8a^2b-2a^4b^4-7)\] and \[(3)(ab)(b^2)= 3ab^3\]